# DiffFuSR: Super-Resolution of all Sentinel-2 Multispectral Bands using Diffusion Models

### v1

@@ Abstract
This paper presents DiffFuSR, a modular pipeline for super-resolving all 12 spectral bands of Sentinel-2 Level-2A imagery to a unified ground sampling distance (GSD) of 2.5 meters. The pipeline comprises two stages: (i) a diffusion-based super-resolution (SR) model trained on high-resolution RGB imagery from the NAIP and WorldStrat datasets, harmonized to simulate Sentinel-2 characteristics; and (ii) a learned fusion network that upscales the remaining multispectral bands using the super-resolved RGB image as a spatial prior. We introduce a robust degradation model and contrastive degradation encoder to support blind SR. Extensive evaluations of the proposed SR pipeline on the OpenSR benchmark demonstrate that the proposed method outperforms current SOTA baselines in terms of reflectance fidelity, spectral consistency, spatial alignment, and hallucination suppression. Furthermore, the fusion network significantly outperforms classical pansharpening approaches, enabling accurate enhancement of Sentinel-2’s 20 m and 60 m bands. This study underscores the power of harmonized learning with generative priors and fusion strategies to create a modular framework for Sentinel-2 SR. Our code and models can be found at https://github.com/NorskRegnesentral/DiffFuSR.
Index Terms:
Super-resolution, Sentinel-2, diffusion models, Wald protocol, data fusion
@@ I Introduction
The democratization of Earth observation has been significantly propelled by the advent of open-access satellite missions such as Sentinel-2, which offer freely available multispectral imagery on a global scale. This data serves as a valuable resource for a broad spectrum of applications, including land cover mapping, agricultural monitoring, urban planning, and disaster response [1]. Despite the relatively high spatial resolution of some Sentinel-2 bands (up to 10 m), a substantial portion of the spectral channels, particularly those at 20 m and 60 m, remain too coarse for tasks that demand fine-grained spatial details. This disparity in resolution across bands limits the full exploitation of the data, especially in applications requiring pixel-level precision. There exist several other satellite sensors that provide RGB and multispectral imagery with higher resolution than Sentinel-2, e.g., WorldView 2, 3 and Legion from Maxar, Pléiades from Airbus, and PlanetScope from Planet Labs. However, this data is not freely available, and the acquisition costs may be substantial. Given the global coverage and open availability of Sentinel-2’s data, enhancing the spatial resolution through computational means is therefore a highly attractive prospect.
@@
To fully leverage the potential of Sentinel-2’s multispectral imagery, super-resolution (SR) has emerged as a compelling approach for enhancing the spatial details [2, 3, 4, 5, 6]. SR refers to reconstructing high-resolution images from their lower-resolution counterparts, often using data-driven or model-based techniques. Mathematically, SR is regarded as an ill-posed problem as there is a non-uniqueness of the solution. For a given low-resolution (LR) image, there are theoretically an infinite number of high-resolution (HR) images that could have produced that LR image when down-sampled. As the down-sampling process inherently loses high-frequency information (details), reversing this process means trying to infer this lost information, and many different ”guesses” can be consistent with the LR input. Moreover, SR is also criticized for its reliance on learned priors and the risk of generating hallucinated details [7]. However, when applied judiciously, with a clear understanding of its limitations, it can significantly improve both the visual interpretability and the utility of satellite imagery in downstream analytical tasks. In the context of Sentinel-2, which consists of multispectral bands in 10 m, 20 m and 60 m resolution, an opportunity arises to combine SR with data fusion techniques to enhance spatial resolution further. Unlike conventional SR, which relies solely on low-resolution inputs, fusion approaches, such as pansharpening, leverage the higher-resolution spectral bands (e.g., the 10 m RGB bands) as guiding information to reconstruct finer details in the lower-resolution bands (e.g., 20 m and 60 m) [8, 9]. This reduces the ambiguity associated with SR by anchoring the reconstruction in real, higher-resolution measurements.
In this work, we propose DiffFuSR, a two-step pipeline that addresses the SR of Sentinel-2 imagery in a comprehensive manner. Our approach begins by super-resolving the 10 m-resolution RGB bands to an enhanced spatial resolution of 2.5 m. Focusing initially on RGB is particularly strategic: the majority of high-resolution reference data available, whether from commercial satellites or aerial imagery, is limited to RGB channels, rather than the full 12-band Sentinel-2 spectrum. Consequently, prioritizing RGB SR allows for effective learning from widely available datasets while remaining consistent with practical downstream applications.
Building upon the super-resolved 2.5 m RGB imagery, we then develop a data fusion strategy to upscale the remaining spectral bands to a unified 2.5 m spatial resolution. For each individual lower-resolution band, the corresponding 2.5 m RGB signal serves as a high-frequency spatial prior, guiding the fusion process. This design allows us to simultaneously preserve the spectral integrity of each band while enriching it with the fine-grained structural information captured in the RGB reference, thus mitigating the risks typically associated with hallucination in purely SR-based approaches.
@@
For the RGB SR step, we develop a diffusion-based model to enhance the spatial resolution of the Sentinel-2 RGB bands to 2.5 m. While diffusion models have gained traction in natural image generation [10, 11, 12], their application to remote sensing SR remains limited. In addition to diffusion models, we leverage large and diverse training datasets and carefully address domain-specific challenges. We utilize a diverse corpus from aerial NAIP imagery [3] and satellite-based WorldStrat data [13], despite their substantial spectral and radiometric differences from Sentinel-2. To bridge these gaps, we apply spectral harmonization techniques [3], which prove critical for generalization. Additionally, we incorporate blind kernel modeling [14] to simulate unknown degradations typical in real-world Sentinel-2 data. Evaluations using the OpenSR benchmark [15] and qualitative results show that our models yield significant performance improvements over current state-of-the-art baselines [16, 17, 18, 3].
@@
For enhancing the resolution of the remaining multispectral bands, we design three dedicated fusion pipelines targeting the 10 m, 20 m, and 60 m bands, respectively. Each fusion model leverages the super-resolved 2.5 m RGB imagery as a high-frequency spatial prior, adopting a structure inspired by pansharpening principles. To enable supervised learning without external HR references, we simulate LR training pairs using the Wald protocol [19] on native Sentinel-2 data. The input configuration is tailored to each task: 10 m bands use only native 10 m inputs; 20 m bands combine 10 m and 20 m data; and 60 m bands integrate 10 m, 20 m, and 60 m channels. This targeted design ensures that each fusion task fully exploits the available spatial information. Evaluations show that our learned fusion models significantly outperform classical methods such as Gram-Schmidt pansharpening [20] in terms of quantitative metrics such as ERGAS, PSNR, and SSIM, as well as improved visual fidelity and spectral consistency.
In summary, this work presents a unified framework for super-resolving all Sentinel-2 bands to a common 2.5 m resolution. Unlike previous methods that have addressed either RGB SR or limited multispectral fusion separately, our approach integrates both into a modular, end-to-end pipeline capable of blind SR across all spatial resolutions and spectral channels. The key contributions are summarized as follows:
@@
  • A novel diffusion-based SR framework for Sentinel-2 RGB bands, leveraging large-scale aerial and satellite datasets, enhanced with data harmonization and blind kernel modeling to ensure robust generalization across diverse acquisition conditions.
@@
  • A new data fusion framework for multispectral super-resolution that enhances the resolution of all Sentinel-2 bands (10 m, 20 m, and 60 m) to a unified 2.5 m spatial resolution using only native Sentinel-2 data, trained using Wald protocol simulations.
@@
The rest of the paper is organized as follows: Section II reviews existing SR models, with a focus on Earth observation applications. Section III details the methodologies that constitute the SR pipeline. Section IV describes the experiments and results, focusing on well-known benchmark datasets. The discussion is provided in Section VI, and finally, Section VII concludes this work.
@@ II Background and related works
Super-resolution (SR) has a rich history, with early breakthroughs focusing on enhancing the resolution by utilizing known relative displacements in image sequences [21], and single image SR using of sparse representations in terms of coupled dictionaries jointly trained from high- and low-resolution image patch pairs [22]. The introduction of deep learning in computer vision was early adapted for SR. Pioneering efforts that demonstrated the potential of convolutional neural network (CNNs) in enhancing image resolution include the SRCNN approach by Dong et al. [23].
@@
Enhancing the resolution of remote sensing images, particularly from satellites, is extremely valuable, as it significantly enhances the utility of the data. Many satellite sensors (e.g., Landsat, WorldView 2 & 3, SPOT 6, …) have therefore been equipped with a panchromatic band in higher resolution, which offers a more detailed view of features on the ground. Pansharpening, i.e., methods that utilize the panchromatic band to enhance the resolution of simultaneously acquired multispectral data, has consequently been a research topic for decades [24].
@@
However, not all satellite sensors have a panchromatic band; hence, single-image SR has also become a central focus in remote sensing, particularly for enhancing the spatial resolution of Sentinel-2 imagery. Early work, such as Lanaras et al. [8], focused on enhancing the resolution of the RGB bands using CNNs trained with synthetic data. Later methods expanded to multispectral SR through data fusion approaches, combining multiple Sentinel-2 revisits [4] or external high-resolution sources such as PlanetScope [6].
@@
The generative power of generative adversarial networks (GANs) has also been explored for SR. Early works include the SRGAN approach by Ledig et al. [25], which were further enhanced by Wang et al. [26]. The benefit of using generative methods like GANs for SR is the potential of creating photo-realistic images, contrary to non-generative methods, which tend to create too smooth results [25]. After gaining influence in computer vision, GANs were also adopted to enhance the resolution of remote sensing images (see e.g., [27]). Recent advances in generative modeling have motivated the exploration of diffusion models for SR tasks [11, 10]. In the remote sensing domain, Wu et al. [14] applied conditional diffusion models for Sentinel-2 SR, enhancing only the RGB bands by leveraging paired SPOT 6/7 imagery. However, their approach remains constrained to RGB bands, and the reliance on paired HR datasets limits its direct scalability to the full 12-band Sentinel-2 MSI.
@@
Efforts such as the WorldStrat dataset [13] and the OpenSR benchmark [15], have provided valuable resources for training and evaluating remote sensing SR models under real-world conditions. Wolters et al. [7] leveraged NAIP aerial imagery for generative SR, but faced challenges due to spectral and radiometric inconsistencies between NAIP and Sentinel-2, leading to domain gaps that degraded the performance. The SEN2NAIP dataset and accompanying work by Aybar et al. [3] introduced a harmonized degradation model to better align NAIP images with Sentinel-2 characteristics, proving crucial for cross-sensor learning. Donike et al. [2] employed latent diffusion models trained on SEN2NAIP, demonstrating that operating in the latent space can significantly reduce resource consumption. They focus solely on super resolving the 10 m Sentinel-2 bands.
@@
Beyond SR, pansharpening techniques have historically been used to fuse high-resolution spatial information with lower-resolution multispectral data [28, 9, 29]. Classical methods such as Gram-Schmidt pansharpening [20] can be effective but may prove insufficient or introduce spectral distortions when attempting to fuse super-resolved RGB images (acting as a pseudo-panchromatic band) with Sentinel-2’s full range of multispectral bands, especially when targeting a significant resolution enhancement to 2.5 m [29]. Lanaras et al. [8] demonstrated that neural networks can perform such fusion more effectively than classical methods, mitigating some distortions. However, their method was limited to enhancing bands to a 10 m ground sampling distance (GSD).
Figure 1: Overview of the proposed pipeline combining RGB super-resolution and multispectral fusion. The top branch illustrates the training of a denoising diffusion model using synthetic low-resolution RGB images generated through harmonization, blurring, downsampling, and noise addition. The model learns to reconstruct high-resolution RGB images. The bottom branch depicts the fusion process: multispectral Sentinel-2 bands are fused with the super-resolved RGB output via a learned fusion module to produce full-resolution 12-band output.
@@
The modular framework presented in this paper significantly extends previous work by jointly addressing super-resolution of all Sentinel-2 bands to a unified 2.5 m resolution. First, we super-resolve the RGB bands using an efficient diffusion model trained on harmonized data, ensuring robust generalization. Second, a dedicated learned fusion network, inspired by pansharpening principles but trained for superior spectral fidelity, upscales the remaining 10 m, 20 m, and 60 m Sentinel-2 bands. While prior works address isolated components of this problem, to the best of our knowledge, this is the first unified approach that performs blind SR across all Sentinel-2 bands and resolutions.
@@ III Method
The goal of this work is to super-resolve all 12 Sentinel-2 multispectral bands to a unified 2.5 m GSD. First in section III-A, we describe a conditional denoising diffusion model to super-resolve the RGB bands to 2.5 m resolution, trained on harmonized synthetic data to ensure robust generalization. Secondly in section III-B, we introduce a learned multispectral fusion network that leverages the super-resolved RGB output to upscale the remaining Sentinel-2 bands to the target resolution of 2.5 m. The overall pipeline is illustrated in Fig. 1. Next, we will describe each module of the pipeline in detail.
@@ III-A Super-Resolution of RGB bands using Conditional Denoising Diffusion Models
We formulate the RGB SR task as conditional generation using a denoising diffusion probabilistic model (DDPM) inspired from [14]. The goal is to generate a HR RGB image $\mathbf{x}_{0}\in\mathbb{R}^{C\times H\times W}$, conditioned on a LR observation $\mathbf{x}_{\text{LR}}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$, where $H^{\prime}<H$ and $W^{\prime}<W$. The model learns to reverse a Gaussian corruption process applied to $\mathbf{x}_{0}$, using the fixed $\mathbf{x}_{\text{LR}}$ to create multiple conditioning inputs. The diffusion model based SR pipeline consists of a forward diffusion process and a reverse denoising process (Fig. 2).
Figure 2: Architecture of the proposed conditional denoising diffusion model for RGB SR. The model learns to reconstruct a high-resolution RGB image $\mathbf{x}_{0}$ from a noisy latent $\mathbf{x}_{T}$ via a denoising process modeled as a Markov chain. The input image is first degraded with Gaussian noise and spatially rearranged using a pixel-folding operation. A spatial encoder extracts contextual features $\mathbf{u}$ from the low-resolution input, while a degradation encoder produces a global degradation embedding $\mathbf{v}$. Both $\mathbf{u}$ and $\mathbf{v}$ are used to condition the denoising network at each reverse time step $t$. The sequence $\mathbf{x}_{T}\rightarrow\dots\rightarrow\mathbf{x}_{0}$ is denoised through a learned reverse diffusion process. The final output is reshaped using a pixel shuffle operation to recover the super-resolved RGB image.

### v2

@@ Abstract
This paper presents DiffFuSR, a modular pipeline for super-resolving all 12 spectral bands of Sentinel-2 Level-2A imagery to a unified ground sampling distance (GSD) of 2.5 meters. The pipeline comprises two stages: (i) a diffusion-based super-resolution (SR) model trained on high-resolution RGB imagery from the NAIP and WorldStrat datasets, harmonized to simulate Sentinel-2 characteristics; and (ii) a learned fusion network that upscales the remaining multispectral bands using the super-resolved RGB image as a spatial prior. We introduce a robust degradation model and contrastive degradation encoder to support blind SR. Extensive evaluations of the proposed SR pipeline on the OpenSR benchmark demonstrate that the proposed method outperforms current SOTA baselines in terms of reflectance fidelity, spectral consistency, spatial alignment, and hallucination suppression. Furthermore, the fusion network significantly outperforms classical and learned pansharpening approaches, enabling accurate enhancement of Sentinel-2’s 20 m and 60 m bands. This work proposes a novel modular framework Sentinel-2 SR that utilizes harmonized learning with diffusion models and fusion strategies. Our code and models can be found at https://github.com/NorskRegnesentral/DiffFuSR.
@@ I Introduction
The democratization of Earth observation has been significantly propelled by the advent of open-access satellite missions such as Sentinel-2, which offer freely available multispectral imagery on a global scale. This data serves as a valuable resource for a broad spectrum of applications, including land cover mapping, agricultural monitoring, urban planning, and disaster response [kansakar2016review]. There exist several other satellite sensors that provide RGB and multispectral imagery with higher resolution than Sentinel-2, e.g., WorldView 2, 3 and Legion from Maxar, Pléiades from Airbus, and PlanetScope from Planet Labs. However, this data is not freely available, and the acquisition costs may be substantial. Given the global coverage and open availability of Sentinel-2’s data, enhancing the spatial resolution through computational means is therefore a highly attractive prospect [donike2025trustworthy, aybar2024sen2naip, 10238735, kowaleczko2023real, rs12152366].
@@
Super resolution (SR) refers to reconstructing high-resolution (HR) images from their lower-resolution counterparts, often using data-driven or model-based techniques. SR is an inherently ill-posed problem, as multiple HR images can produce the same low-resolution (LR) counterpart, requiring models to infer lost high-frequency details. Moreover, SR is also criticized for its reliance on learned priors and the risk of generating hallucinated details [satlassuperres]. However, when applied judiciously, with a clear understanding of its limitations, it can significantly improve both the visual interpretability and the utility of satellite imagery in downstream analytical tasks.
In the context of Sentinel-2, which consists of multispectral bands in 10 m, 20 m and 60 m resolution, an opportunity arises to combine SR with data fusion techniques to enhance spatial resolution further. Unlike conventional SR, which relies solely on low-resolution inputs, fusion approaches, such as pansharpening, leverage the higher-resolution spectral bands (e.g., the 10 m RGB bands) as guiding information to reconstruct finer details in the lower-resolution bands  [lanaras2018super, vivone2014critical]. This reduces the ambiguity associated with SR by anchoring the reconstruction in real, higher-resolution measurements.
While diffusion models have gained traction in natural image SR [rombach2022high, saharia2022image, teng2023relay], their application to remote sensing SR remains limited. In this work, we propose DiffFuSR, a two-step pipeline that begins by super-resolving Sentinel-2’s 10 m RGB bands to 2.5 m using a diffusion model trained on harmonized aerial and Sentinel-2 data. Subsequently, a learned fusion network upscales the remaining spectral bands, using the super-resolved RGB as a spatial prior to ensure spectral integrity and mitigate hallucination.
@@
We leverage large and diverse training datasets and carefully address domain-specific challenges. We utilize a diverse corpus from aerial NAIP imagery [aybar2024sen2naip] and satellite-based WorldStrat data [cornebise2022open], despite their substantial spectral and radiometric differences from Sentinel-2. To bridge these gaps, we apply spectral harmonization techniques [aybar2024sen2naip], which prove critical for generalization. Additionally, we incorporate blind kernel modeling [wu2023conditional] to simulate unknown degradations typical in real-world Sentinel-2 data. Evaluations using the OpenSR benchmark [opensrtest] and qualitative results show that our models yield significant performance improvements over current state-of-the-art baselines [wolters2023zooming, cresson2022sr4rs, 10887321, aybar2024sen2naip].
@@
To enhance the resolution of the remaining multispectral bands, we design a neural network-based fusion module and train three instances of it for 10 m, 20 m, and 60 m bands, respectively. Each fusion model leverages the super-resolved 2.5 m RGB imagery as a high-frequency spatial prior, adopting a structure inspired by pansharpening principles. To enable supervised learning without external HR references, we simulate LR training pairs using the Wald protocol [wald2002data] on native Sentinel-2 data. This design ensures that the fusion task fully exploits the available spatial information. Evaluations show that our learned fusion models significantly outperform classical and learned baseline methods in terms of quantitative metrics such as ERGAS, PSNR, and SSIM, as well as improved visual fidelity and spectral consistency.
In summary, DiffFuSR is a unified framework for super-resolving all Sentinel-2 bands to a common 2.5 m resolution. The key contributions are summarized as follows:
@@
  • A novel diffusion-based SR framework for Sentinel-2 RGB bands, leveraging large-scale aerial and satellite datasets, enhanced with data harmonization and blind kernel modeling.
@@
  • A new data fusion framework for multispectral SR that enhances the resolution of all Sentinel-2 bands (10 m, 20 m, and 60 m) to a unified 2.5 m spatial resolution using only native Sentinel-2 data, trained using Wald protocol simulations.
@@ II Background and related works
Early SR methods evolved from multi-image registration [irani1991improving] and sparse representations [yang2010image] to pioneering deep learning approaches like SRCNN [dong2015image], which established convolutional neural networks (CNNs) as the standard for the task.
@@
Pansharpening, i.e., methods that utilize the panchromatic band available in many satellites (e.g., Landsat, WorldView etc.) to enhance the resolution of simultaneously acquired multispectral data, has been a research topic for decades [deng2022machine].
@@
However, not all satellite sensors have a panchromatic band; hence, single-image SR has also become a central focus in remote sensing, particularly for enhancing the spatial resolution of Sentinel-2 imagery. Early work, such as Lanaras et al. [lanaras2018super], focused on enhancing the resolution of the RGB bands using CNNs trained with synthetic data. Later methods expanded to multispectral SR through data fusion approaches, combining multiple Sentinel-2 revisits [10238735] or external HR sources such as PlanetScope [rs12152366].
@@
The generative power of generative adversarial networks (GANs) has also been explored for SR. Early works include the SRGAN approach by Ledig et al. [ledig2017photo], which were further improved by Wang et al. [wang2018esrgan]. The benefit of using generative methods like GANs for SR is the potential of creating photo-realistic images, contrary to non-generative methods, which tend to create too smooth results [ledig2017photo]. Recent advances in generative modeling have motivated the exploration of diffusion models for SR tasks [saharia2022image, rombach2022high]. Wu et al. [wu2023conditional] applied conditional diffusion models for Sentinel-2 SR, enhancing only the RGB bands by leveraging paired SPOT 6/7 imagery. However, their approach remains constrained to RGB bands, and the reliance on paired HR datasets limits its direct scalability to the remaining Sentinel-2 bands.
@@
Efforts such as the WorldStrat dataset [cornebise2022open] and the OpenSR benchmark [opensrtest], have provided valuable resources for training and evaluating remote sensing SR models under real-world conditions. Wolters et al. [satlassuperres] leveraged NAIP aerial imagery for generative SR, but faced challenges due to spectral and radiometric inconsistencies between NAIP and Sentinel-2, leading to domain gaps that degraded the performance. The SEN2NAIP dataset and accompanying work by Aybar et al. [aybar2024sen2naip] introduced a harmonized degradation model to better align NAIP images with Sentinel-2 characteristics, proving crucial for cross-sensor learning. Donike et al. [donike2025trustworthy] employed latent diffusion models trained on SEN2NAIP, demonstrating that operating in the latent space can significantly reduce resource consumption. They focus solely on super-resolving the 10 m Sentinel-2 bands.
@@
After performing SR on the 10 m bands, pansharpening techniques can be applied to fuse the super-resolved image with the lower-resolution multispectral bands (e.g., 20 m and 60 m) [sarmad2024diffusion]. Classical methods such as Gram-Schmidt pansharpening [dalla2015global] can be effective but may prove insufficient or introduce spectral distortions when attempting to fuse super-resolved RGB images (acting as a pseudo-panchromatic band) with Sentinel-2’s remaining multispectral bands, especially when targeting a significant resolution enhancement to 2.5 m [sarmad2024diffusion]. Lanaras et al. [lanaras2018super] demonstrated that neural networks can perform such fusion more effectively than classical methods, mitigating some distortions. However, their method was limited to enhancing bands to a 10 m ground sampling distance (GSD) instead of a higher resolution of 2.5 m GSD.
@@
The modular framework presented in this paper significantly extends previous work by jointly addressing SR of all Sentinel-2 bands to a unified 2.5 m resolution. First, we super-resolve the RGB bands using an efficient diffusion model trained on harmonized data, ensuring robust generalization. Second, a learned fusion network, inspired by pansharpening principles but trained for superior spectral fidelity, upscales the remaining 10 m, 20 m, and 60 m Sentinel-2 bands. While prior works address isolated components of this problem, to the best of our knowledge, this is the first unified approach that performs blind SR across all Sentinel-2 bands and resolutions.
@@ III Method
First, in section III-A, we describe a conditional denoising diffusion model to super-resolve the RGB bands to 2.5 m GSD, trained on harmonized synthetic data to ensure robust generalization. Secondly, in section III-B, we introduce a learned multispectral fusion network that leverages the super-resolved RGB output to upscale the remaining Sentinel-2 bands to the target GSD of 2.5 m. The overall pipeline is illustrated in Fig. 1.
Figure 1: Overview of the proposed pipeline combining RGB SR and multispectral fusion. The top branch illustrates the training of a denoising diffusion model using synthetic low-resolution RGB images generated through harmonization, blurring, downsampling, and noise addition. The model learns to reconstruct HR RGB images. The bottom branch depicts the fusion process: multispectral Sentinel-2 bands are fused with the super-resolved RGB output via a learned fusion module to produce full-resolution 12-band output.
@@ III-A Super-Resolution of RGB bands using Conditional Denoising Diffusion Models
We formulate the RGB SR task as conditional generation using a denoising diffusion probabilistic model (DDPM) inspired from [wu2023conditional]. The goal is to generate a HR RGB image $\mathbf{x}_{0}\in\mathbb{R}^{C\times H\times W}$, conditioned on a LR observation $\mathbf{x}_{\text{LR}}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$, where $H^{\prime}<H$ and $W^{\prime}<W$. The model learns to reverse a Gaussian noise corruption process added to $\mathbf{x}_{0}$, using the fixed $\mathbf{x}_{\text{LR}}$ to create multiple conditioning inputs. The SR pipeline consists of a forward diffusion process and a reverse denoising process (Fig. 2).
Figure 2: Architecture of the proposed conditional denoising diffusion model for RGB SR. The model learns to reconstruct a HR RGB image $\mathbf{x}_{0}$ from a noisy latent $\mathbf{x}_{T}$ via a denoising process modeled as a Markov chain. The input image is first degraded with Gaussian noise and spatially rearranged using a pixel-folding operation. A spatial encoder extracts contextual features $\mathbf{u}$ from the low-resolution input, while a degradation encoder produces a global degradation embedding $\mathbf{v}$. Both $\mathbf{u}$ and $\mathbf{v}$ are used to condition the denoising network at each reverse time step $t$. The sequence $\mathbf{x}_{T}\rightarrow\dots\rightarrow\mathbf{x}_{0}$ is denoised through a learned reverse diffusion process. The final output is reshaped using a pixel shuffle operation to recover the super-resolved RGB image.